Determine the numeric value of this postfix expression
{ 4, 2, -, 3, 6, /, 2, *, / }

Answer by Jk22(389)

{ 4, 2, -, 3, 6, /, 2, *, / }
4-2 = 2
3/6 = 1/2
1/2*2 = 1
2/1=2